OpenStudy (needhelpinalgebra):

Can somebody help me out with an algebra question? Will Fan and Medal.

OpenStudy (writerforever18):

Whats your question?

OpenStudy (needhelpinalgebra):

The graph plots four equations, A, B, C, and D: Which pair of equations has (0, 8) as its solution? (4 points) Equation A and Equation C Equation B and Equation C Equation C and Equation D Equation B and Equation D

OpenStudy (liz_beauty45):

wow so may owls

OpenStudy (liz_beauty45):

many

OpenStudy (needhelpinalgebra):

lol

OpenStudy (liz_beauty45):

i remember when i was 1st a owl

OpenStudy (needhelpinalgebra):

can you help me with this, @Liz_Beauty45 , @writerforever18

OpenStudy (liz_beauty45):

i would love to....but i suck at math

OpenStudy (gamer_girl2413):

its A and c

OpenStudy (needhelpinalgebra):

Same. I am more of a science nerd. Hi, @gamer_girl2413

OpenStudy (liz_beauty45):

yeah i LOVE science

OpenStudy (gamer_girl2413):

hey

OpenStudy (writerforever18):

Sorry I just came back

OpenStudy (needhelpinalgebra):

LMAO could you help out with a few more? there are close to 22 more..... practice test. :(

OpenStudy (gamer_girl2413):

sure

OpenStudy (needhelpinalgebra):

A pair of linear equations is shown below: y = −x + 1 y = 2x + 4 Which of the following statements best explains the steps to solve the pair of equations graphically? (4 points) On a graph, plot the line y = −x + 1, which has y-intercept = −1 and slope = 1, and y = 2x + 4, which has y-intercept = 2 and slope = 4, and write the coordinates of the point of intersection of the two lines as the solution. On a graph, plot the line y = −x + 1, which has y-intercept = 1 and slope = 1, and y = 2x + 4, which has y-intercept = 1 and slope = 4, and write the coordinates of the point of intersection of the two lines as the solution. On a graph, plot the line y = −x + 1, which has y-intercept = 1 and slope = −1, and y = 2x + 4, which has y-intercept = −2 and slope = 2, and write the coordinates of the point of intersection of the two lines as the solution. On a graph, plot the line y = −x + 1, which has y-intercept = 1 and slope = −1, and y = 2x + 4, which has y-intercept = 4 and slope = 2, and write the coordinates of the point of intersection of the two lines as the solution.

OpenStudy (gamer_girl2413):

idk sorry but i can try to look it up

OpenStudy (needhelpinalgebra):

that is what im doing now. lmao

OpenStudy (gamer_girl2413):

yeah i didn't find anything u

OpenStudy (needhelpinalgebra):

still looking.

OpenStudy (needhelpinalgebra):

found it.

OpenStudy (needhelpinalgebra):

apparently the last one.

OpenStudy (needhelpinalgebra):

The two lines, A and B, are graphed below: Determine the solution and the reasoning that justifies the solution to the systems of equations. (4 points) (−4, 6), because both the equations are true for this point (2, 8), because the graph of the two equations intersects at this point (2, 8), because neither of the two equations are true for this coordinate point (−4, 6), because the graph of the two equations intersects the x-axis at these points

OpenStudy (needhelpinalgebra):

I believe it is 2,8 because the lines intersect at that point. It could be the other one, though...

OpenStudy (gamer_girl2413):

I'm pretty sure it b

OpenStudy (needhelpinalgebra):

same.

OpenStudy (needhelpinalgebra):

A system of equations is shown below: 6x − 2y = 3 (equation 1) 5x + 3y = 4 (equation 2) A student wants to prove that if equation 2 is kept unchanged and equation 1 is replaced with the sum of equation 1 and a multiple of equation 2, the solution to the new system of equations is the same as the solution to the original system of equations. If equation 2 is multiplied by 1, which of the following steps should the student use for the proof? (4 points) Show that the solution to the system of equations 10x + y = 7 and 5x + 3y = 4 is the same as the solution to the given system of equations Show that the solution to the system of equations 10x − 2y = 7 and 5x + 3y = 4 is the same as the solution to the given system of equations Show that the solution to the system of equations 11x − y = 7 and 5x + 3y = 4 is the same as the solution to the given system of equations Show that the solution to the system of equations 11x + y = 7 and 5x + 3y = 4 is the same as the solution to the given system of equations

OpenStudy (gamer_girl2413):

give me a sec I'm working on y module one test part 2

OpenStudy (gamer_girl2413):

my*

OpenStudy (needhelpinalgebra):

the written part?

OpenStudy (gamer_girl2413):

yeah

OpenStudy (needhelpinalgebra):

lol

OpenStudy (gamer_girl2413):

do you know the answer to q 2 part b?

OpenStudy (needhelpinalgebra):

maybe. what is the question?

OpenStudy (gamer_girl2413):

Using the information in Part A, interpret the meaning of the quotient in terms of the two fractions given. 2/5 4 and 4/5

OpenStudy (needhelpinalgebra):

hmm. I just look it up on google, modify the answers found on there.

OpenStudy (gamer_girl2413):

okay

OpenStudy (needhelpinalgebra):

lol. got another one....

OpenStudy (needhelpinalgebra):

Two systems of equations are shown below: System A System B 6x + y = 2 2x − 3y = −10 −x −y = −3 −x − y = −3

OpenStudy (needhelpinalgebra):

The value of x for System B will be 4 less than the value of x for System A because the coefficient of x in the first equation of System B is 4 less than the coefficient of x in the first equation of System A. They will have the same solution because the first equations of both the systems have the same graph. They will have the same solution because the first equation of System B is obtained by adding the first equation of System A to 4 times the second equation of System A. The value of x for System A will be equal to the value of y for System B because the first equation of System B is obtained by adding −4 to the first equation of System A and the second equations are identical.

OpenStudy (needhelpinalgebra):

@AloneS

OpenStudy (needhelpinalgebra):

its all the way down, @AloneS

alones (alones):

ummm.. can you make a new post?

OpenStudy (needhelpinalgebra):

yup. got new questions anyways. XD sowwy

alones (alones):

lool np hun ;)